Session: Feb 2011 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Pablo de Vicente at Sat Mar 12 07:42:56 2011) Success: HIGH FS
Field System froze between 07:05 and 08:56. However the data were not good until 09:19 Very bad weather causing high Tsys since Friday evening (last 12 hours of the experiment)

Onsala (Michael Lindqvist at Sat Mar 12 08:15:02 2011) Success:
Murphy's law: "Anything that can go wrong, will go wrong". A lot of problems between approx 22-08 UTC (March 10-11). * Network problems to JIVE * Mark5 problems. * To high winds. We had to stow the telescope * FS-computer hard disk crashed. Third times within a month!!! Luckily we have backups. * Short power outage which affected the telescope, VLBI-rack,.... * ....

Medicina (Zacchiroli at Sat Mar 12 10:00:28 2011) Minor failures: RFI

Westerbork (Geert Kuper at Sat Mar 12 10:32:19 2011) Success:
No known problems

Torun (Kaz Borkowski at Mon Mar 14 08:15:28 2011) Success:
No known problems.

Effelsberg (Uwe Bach at Thu Mar 17 13:23:31 2011) Success:

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:35:25 2011) Minor failures: WIND
Telescope was parked from 12:48 to 14:14 and from 15:48 to 17:07 (day 69) due to high winds. Some instances of 2-3 minutes lost due to antenna control problems. No other known problems.